OpenStudy (anonymous):

Can someone please help me? Step by step please. A car is coasting backwards downhill at a speed of 3.0m/s when the driver gets the engine started. After 2.5 s the car is moving uphill at 4.5m/s. If uphill is chosen as the positive direction, what is the car's average acceleration?

OpenStudy (anonymous):

Let's do the GUESS method WEEEE, so you know your givens are initial velocity (3.0m/s), final velocity (4.5 m/s) and the time it took to reach final velocity (2.55 seconds) G: Vi= 3.0 m/s Vf= 4.5 m/s T= 2.55 seconds Now they ask to find the average acceleration so your unknowns just look like.. U: Average Acceleration= ? E: Avg acceleration= (Vf-Vi)/ change in time (equation to find avg acceleration) now you substitute and note that the positive direction is uphill and downhill is the negative direction, so that means 3.0 m/s is negative and 4.5 m/s is positive S: \[Aavg= 4.5 m/s - (-3.0 m/s) \div 2.55\] S: 2.94= Avg. acceleration